Can u let me if any screw compressor discharge pressure is 2 bar then how much compressor capacity approximately.
1
Expert's answer
2021-08-27T01:38:01-0400

compressor capacity :

"A = Q.\frac{\beta}{\eta}="

"Q =238.48 \frac{l}{min}"

"\beta = 1.3"

"\eta = 0.75"

"A = Q.\frac{\beta}{\eta}= 283.48 \frac{1.3}{0.75}= 491.36 \frac{l}{min}"

where

Q – total air volume to be consumed by all consumers of pneumatic system to be measured in l/min;

β – coefficient to count for design specifics of compressor unit by manufacturer;

η – performance index of compressor unit.